How is the 38th percentile wage calculated?
Answer #1

Solution:

The 38th percentile refers to a value such that at least 38% of the observations is equal to or less than this value and at least 62% (=100-38 = 62) of the observations is equal to or greater than this value.

The first step will be to sort the given values

Compute n i.e. number of values

Compute the index i : 38/100 * n = 0.38n

If the index i is an integer, then 38th percentile will be average of the values at the positions i and i+1

It will be the 38th percentile value